Given dy/dx = y-x,where y(0) = 2, find y(0.1) and y(0.2) by Euler’s method up to two decimal places.


1
Expert's answer
2021-02-01T07:34:17-0500

Solution

Euler method is a numerical procedure for solving ordinary differential equations with a given initial value.

For

y’(x) = f(x,y(x)), y(x0)=y0 

the size of every step h and xn = x0+n*h an approximation of the solution to the ODE is

yn+1 = yn + h*f(xn,yn)  

In this case x0 = 0, y0 = 2, f(x,y) = y – x, h = 0.1

So

yn+1 = yn + h*(yn – xn ) = (1 + h)*yn – h*xn  

y1 = y(0.1) = 2.2,  y2 = y(0.2) = 2.41

Answer

y(0.1) = 2.2,  y(0.2) = 2.41